OpenStudy (anonymous):

HELP PLEASE!! solve by substitution method. 4x+5y=2 x=10-6y

OpenStudy (trancenova):

As x=10-6y put x into 4x+5y=2 4(10-6y)+5y=2 expand 40-24y+5y=2 -19y=-38 y=2 Put y = 2 into x=10-6y to find x